During an adiabatic process, the temperature of 6.
Answer #1

Work done adiabatic process is given as

W=\frac{P_1V_1-P_2V_2}{\gamma -1}

Not at 1st and 2nd state

P_1V_1=nRT_1\\ P_2V_2=nRT_2

hence

\Delta W=\frac{nR(T_1-T_2)}{2.5 -1}=\frac{6.10\times8.31\times(495-147)}{1.5}=1.176\times10^4$ J$

(b) Since the process is adiabatic hence heat exchange will be 0 i.e.

\Delta Q=0

(c)

From 1st law of thermodynam,ics

\Delta U=\Delta Q+\Delta W=0+1.176\times10^4=1.176\times10^4